OpenStudy (anonymous):

simplify sqrt7x+3+2=8

OpenStudy (saifoo.khan):

7x + 3 + 2 = 64

OpenStudy (anonymous):

x=9/7

OpenStudy (saifoo.khan):

It totally depends on what you wrote.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

squaring on both sides

OpenStudy (anonymous):

asuming sqrt(7*x)

OpenStudy (anonymous):

i got x=59/7

OpenStudy (anonymous):

no, its 9/7th

OpenStudy (anonymous):

you are right jenni

OpenStudy (lalaly):

is it sqrt(7) x or sqrt (7x)

OpenStudy (anonymous):

how did you get it

OpenStudy (anonymous):

x =59/7 by squaring both sides

OpenStudy (anonymous):

Jenni, it's not clear what the question is - what is actually inside the square root ?

OpenStudy (anonymous):

thats what i got here let me show my work

OpenStudy (anonymous):

Star, you are wrong...

OpenStudy (anonymous):

Is it this ? \[\sqrt{7x+3}+2=8\]

OpenStudy (anonymous):

@fiddlearound yes

OpenStudy (anonymous):

then you are wrong

OpenStudy (anonymous):

ah, question was mistyped x = 33/7

OpenStudy (anonymous):

so, 7x+3=36

OpenStudy (saifoo.khan):

7x+ 3 = 36 Solve.

OpenStudy (saifoo.khan):

7x = 33 x = 33/7

OpenStudy (anonymous):

so i just subtract the 2 first

OpenStudy (saifoo.khan):

Right

OpenStudy (anonymous):

That's what the radical pros all do Jenni!

OpenStudy (anonymous):

but then it would be x= 33/7

OpenStudy (anonymous):

that's right.

OpenStudy (anonymous):

becasuee 7x=33 i need x alone so divide both by 7

OpenStudy (saifoo.khan):

right